Using laws of exponents, simplify and write the answer in exponential form:  2^ 5 × 5 ^5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of multiplying the powers with same exponents = a^m× b^m= ab^m
Then,
= (2 × 5)^5
= 10^5

Thanks